# MSCPT: Few-shot Whole Slide Image Classification with Multi-scale and Context-focused Prompt Tuning

### v1

@@ IV-B Comparisons with State-of-the-Art
GCN+KNN(Coord.) 1.35M 92.85±2.43 79.29±3.98 81.63±3.74
GCN+KNN(Feat.) 1.35M 93.92±2.66 80.46±4.33 82.41±4.26
GAT+Sim. 1.35M 93.14±1.78 80.82±4.25 82.41±3.85
GraphSAGE+Sim. 2.40M 93.60±2.72 80.40±3.89 82.66±3.69
GCN+Sim.(ours) 1.35M 95.04±1.31 82.59±2.14 85.62±2.14
Prompt tuning methods explicitly designed for WSI exhibit superior performance. This is attributed to their incorporation of priors into pre-trained VLMs and leveraging those priors to guide prompt tuning. Additionally, ViLa-MIL introduces multi-scale information compared to TOP, positioning it as the second-best overall performer. In comparison to ViLa-MIL, MSCPT shows improvements across all datasets, with AUC increasing by 0.9-5.7%, F1 by 2.2-6.2%, and ACC by 2.3-5.5%. This is because MSCPT performs prompt tuning both on the textual and visual modality. Furthermore, MSCPT takes into account both the multi-scale and contextual information of WSIs. Unlike the late fusion approach in ViLa-MIL, MSCPT employs an intermediate fusion method, leveraging the transformer layer and trainable global prompts to integrate pathological visual descriptions from both high and low levels.
IV-C Ablation Experiment
IV-C1 Effects of Each Component in MSCPT
@@
To verify the effectiveness of three core components, ablation experiments were conducted on the TCGA-NSCLC dataset based on PLIP, the experimental results are presented in Table II. When all modules were removed, MSCPT regresses to the baseline (i.e., Metaprompt). All metrics showed significant improvement (2.2%-2.9%) after adding the Multi-scale Hierarchical Prompt Tuning (MHPT) module to the baseline. This is because the MHPT module utilizes transformer layers to integrate pathological visual descriptions across different scales, enhancing the model’s information aggregation capabilities. Building upon this, we introduced the Image-text Similarity-based Graph Prompt Tuning (ISGPT) module, which led to improvements in all metrics (2.4%-2.7%). This demonstrated that utilizing ISGPT for contextual learning also enhances model performance, reaffirming the importance of contextual information for WSI analysis. When we added both the MHPT and Non-Parametric Cross-Guided Pooling (NPCGP) module to the baseline, in comparison to solely adding MHPT, the metrics improved by 1.7%-1.8%. This indicates that the NPCGP module, compared to attention-based pooling, is more effective in identifying important patches within the WSI, resulting in better instance aggregation results. When all modules work together, the baseline is transformed into MSCPT. MSCPT has shown improvements of 7.6% in AUC, 6.1% in F1, and 6.5% in ACC compared to the baseline.
TABLE IV: Ablation experiment of different instance aggregation methods on the TCGA-NSCLC dataset based on PLIP.
Methods TCGA-NSCLC (PLIP-based)
@@
TABLE V: Results of different large language models on the TCGA-RCC dataset based on CLIP.
Methods TCGA-RCC (CLIP-based)
@@
Figure 4: Visualization of the original WSI, similarity score map for patch selection, heatmaps generated using MSCPT, attention heatmaps of the baseline (i.e., Metaprompt) and the best-performing traditional MIL-based method, selected high similarity patches and patches with highest similarity scores using MSCPT. The area surrounded by the red line in the original WSI is the tumor area.
Figure 5: Experiments on TCGA-BRCA with 16, 8, 4, 2, and 1-shot settings. (a) (b) are CLIP-based results, while (c) (d) are PLIP-based results.
IV-C2 Effects of Graph Construction
To validate the effectiveness of building adjacency matrices based on image-text similarity, we used K-Nearest Neighbor (KNN) to create adjacency matrices based on patch coordinates or visual features as referenced in studies [29, 30]. Additionally, we tested the effectiveness of GCN by comparing them with GAT [46] and GraphSAGE [47]. Experimental results on TCGA-RCC using CLIP are shown in Table III, where KNN(Coord.) and KNN(Feat.) refer to using KNN for constructing adjacency matrices based on patch coordinates and patch features, and Sim. signifies constructing adjacency matrices using image-text similarity. Switching to KNN to construct the adjacency matrix led to decreased performance across all metrics, whether based on coordinates or patch visual features. This decline may be attributed to the limited scope of connectivity in the adjacent matrix construction methods. Connecting only nearby patches based on coordinates restricts the GNN to the local context, while connecting visually similar patches based on their features may lack global information about interactions between different types of tissue organization. In contrast, our ISGPT module connects patches related to specific cancer types, overcoming local or visual similarity connection limitations and enabling a more comprehensive contextual understanding. When we replaced GCN with GAT or GraphSAGE, the model’s performance also experienced varying degrees of decline. We believe complex graph neural networks are unsuitable for few-shot scenarios.
IV-C3 Effects of Instance Aggregation
To validate the effectiveness of our instance aggregation method, we compared Non-Parametric Cross-Guided Pooling (NPCGP) with other aggregation methods (i.e., Mean Pooling, Max Pooling, and Attention-based Pooling). The experimental results based on PLIP on the TCGA-NSCLC are presented in Table IV. When we replaced NPCGP with other methods, the performance of the models decreased to varying degrees. This implies that our NPCGP can discern more impactful patches and aggregate them into bag features. The visualization results in Section IV-D support this point. We also conducted an ablation study on cross-guidance. Instead of computing cross-scale cosine similarity during feature aggregation, we only calculated cosine similarity between patch and description embeddings at the same scale. Removing cross-guidance led to a drop in performance across all metrics, likely because LLMs may produce descriptions with incorrect scales.
IV-C4 Effects of Large Language Models
@@
To verify the impact of different LLMs on model performance, we compared the performance of MSCPT when using descriptions generated by different LLMs (i.e., Gemini-1.5-pro [42], Cluade-3 [43], Llama-3 [44], GPT-3.5 [45] and GPT-4 [41]). The results obtained using CLIP on TCGA-RCC are presented in Table V. When generating descriptions using Claude-3, MSCPT performs comparably to the baseline. However, MSCPT outperforms the baseline when using other LLMs. This demonstrates MSCPT’s robustness across different LLMs and highlights the benefit of accurate pathological visual descriptions for model performance. Providing a more accurate description helps improve model performance.
IV-D Visualization
@@
As shown in Fig. 4, we have visualized a case of TCGA-RCC based on PLIP and a case of TCGA-BRCA based on CLIP. As depicted in Fig. 4a, during patch selection, CLIP assigned high similarity scores not only to tumor regions but also to non-tumor areas. This outcome arose because CLIP was not specifically designed for pathological images, resulting in a less-than-optimal zero-shot capability for this type of imagery. However, after prompt tuning using MSCPT, the model correctly assigned high scores to the actual tumor regions, while the regions that originally received high scores dropped to lower score ranges (red arrows in Fig. 4a). Meanwhile, CLAM-MB struggled to differentiate tumor and no-tumor. Similarly, Metaprompt assigned high attention weights to certain non-tumor tissues (red arrows in Fig. 4a).
@@
During selecting patches using PLIP, the model could roughly identify tumor regions but also assigned high scores to a small number of non-tumor areas. However, this issue was mitigated with MSCPT (red arrows in Fig. 4b). While ABMIL could also determine instance importance, it tended to assign higher scores to certain non-tumor regions compared to MSCPT (yellow arrows in Fig. 4b). Due to PLIP’s improved ability to represent pathological images, Metaprompt produced visualization results comparable to MSCPT.
IV-E Results with Fewer Training Samples
To further validate MSCPT’s performance, we conducted experiments on TCGA-BRCA with 16, 8, 4, 2, and 1-shot settings. Based on the results in Table I, we selected several well-performing models (i.e., CLAM-MB, CoOp, Metaprompt, ViLa-MIL, and MSCPT) for these experiments. It is also worth noting that with limited training samples, sample selection significantly impacts model performance [8]. To address this, we conducted dataset splitting, model training, and testing using ten different seeds, excluding the two best and two worst results to calculate the average. Due to the sample imbalance in TCGA-BRCA, we just reported AUC and macro F1-score, as shown in Fig. 5. When using CLIP as the base model, MSCPT underperforms compared to CLAM-MB and Metaprompt in 1- and 2-shot settings, likely due to MSCPT’s larger parameter size and CLIP’s limited understanding of pathology descriptions. However, with 4 or more shots, MSCPT significantly outperforms other methods. Additionally, when using PLIP as the base model, MSCPT consistently performs better than any other method.
V Conclusion
In this paper, we propose Multi-Scale and Context-focused Prompt Tuning (MSCPT) to solve the Few-shot Weakly-supervised WSI Classification (FSWC) task. MSCPT generates multi-scale pathological visual descriptions using GPT-4, guiding hierarchical prompt tuning and instance aggregation. Experiments on three WSI subtyping datasets and two Vision-Language models (VLMs) show that MSCPT achieved state-of-the-art results in FSWC tasks. Furthermore, MSCPT is applicable to fine-tune any VLMs for WSI-level tasks. However, we find that the model performance varies significantly on different datasets and VLMs. That is because the performance of model fine-tuning largely depends on the pre-trained VLM itself. We look forward to the emergence of more comprehensive and powerful pre-trained pathological VLMs, which will significantly promote the development of FSWC tasks and even all computational pathology.

### v2

@@ 4.3 Results with Fewer Training Samples
MSCPT w/ TransMIL 2.44M 86.26±4.74 68.54±6.70 72.13±6.87
MSCPT w/ FA 2.45M 91.12±4.14 71.40±5.88 76.94±5.13
GCN w/ KNN(Coord.) 1.35M 92.85±2.43 79.29±3.98 81.63±3.74
GCN w/ KNN(Feat.) 1.35M 93.92±2.66 80.46±4.33 82.41±4.26
GAT w/ Sim. 1.35M 93.14±1.78 80.82±4.25 82.41±3.85
GraphSAGE w/ Sim. 2.40M 93.60±2.72 80.40±3.89 82.66±3.69
GCN w/ Sim.(ours) 1.35M 95.04±1.31 82.59±2.14 85.62±2.14
Figure 4: Ablation experiment of patch selection. (a) Influence of selected patch counts at 5× magnification in visual prompt tuning. (b) Influence of prompt set counts in patch selection.
Moreover, we observed some phenomena that differ from those in the 16-shot setting. Prompt-tuning methods explicitly designed for WSI experienced a significant drop in performance, particularly under the 2-shot setting. This decline can be attributed to both TOP and ViLa-MIL relying on CLIP, which struggles to comprehend complex and detailed pathological prior knowledge. As a result, these methods provide the VLM with only simplistic pathological priors. Such simplistic priors are challenging to fully exploit under extremely limited data conditions, leading to suboptimal performance. In contrast, MSCPT offers the VLM comprehensive and detailed pathological visual priors across multiple scales, enabling it to leverage all available information, even when data is scarce. Metaprompt also demonstrates superior performance under limited training data, thanks to its ability to capture multi-scale information and utilize visual prompt tuning. Our MSCPT effectively integrates these advantages, resulting in robust performance.
4.4 Ablation Experiment
4.4.1 Effects of Each Component in MSCPT
To verify the effectiveness of the five core components, ablation experiments were conducted on the TCGA-NSCLC and PANDA datasets. The experimental results are presented in Table 3. When all modules were removed, MSCPT regressed to CoOp [22]. Integrating a multi-scale framework with visual prompt tuning into CoOp aligns its implementation with that of Metaprompt. The incorporation of visual prompt tuning and multi-scale information enhances the model’s performance, particularly in scenarios with limited training samples. While CoOp and Metaprompt use class names as text prompts, we introduce more detailed visual prior knowledge as their new text prompts. However, experimental results demonstrate that the performance of CoOp and Metaprompt did not consistently improve after incorporating visual prior knowledge; in fact, there was even a slight decrease in performance. We argue that the prompt tuning methods used in CoOp and Metaprompt, which are designed for natural images, fail to fully leverage the structure of prior knowledge. Furthermore, the use of excessively long text prompts negatively impacts the results.
@@ 4.4.1 Effects of Each Component in MSCPT
All metrics improved by 0.8-2.9% after integrating the Multi-scale Hierarchical Prompt Tuning (MHPT) module into Metaprompt. This is because the MHPT module utilizes transformer layers to integrate pathological visual descriptions across different scales, enhances the model’s information aggregation capabilities. Unlike CoOp, Metaprompt, and ViLa-MIL, which can only incorporate a single, relatively long visual prior knowledge prompt, MSCPT inputs a large amount of detailed visual prior knowledge in separate fragments. This enhances MSCPT’s ability to comprehend and leverage prior knowledge more effectively. Building upon this, we introduced the Image-text Similarity-based Graph Prompt Tuning (ISGPT) module, which led to improvements in all metrics (1.2-2.7%). This demonstrated that utilizing ISGPT for contextual learning also enhances model performance, reaffirming the importance of contextual information for WSI analysis. When we added both the MHPT and Non-Parametric Cross-Guided Pooling (NPCGP) module to the baseline, in comparison to solely adding MHPT, the metrics improved by at most 1.8%. This indicates that the NPCGP module, compared to attention-based pooling, is more effective in identifying important patches within the WSI, resulting in better instance aggregation results. Additionally, NPCGP uses image-text similarity for classification, offering more human-interpretable results compared to attention-based methods.
Furthermore, the importance of the multi-scale architecture in MSCPT was validated by removing the low-level encoders. Without low-level descriptions, the MHPT module retained only global and high-level prompts, while ISGPT and NPCGP were simplified to single-scale structures. As shown in Table 3, this resulted in 4.4% and 6.0% decreases in F1-scores on the TCGA-NSCLC and PANDA datasets, respectively. These results show that the multi-scale structure, despite higher computational complexity, significantly improves performance.
4.4.2 Effects of Patch Selection
Unlike previous prompt tuning methods designed for WSIs, MSCPT incorporates visual prompt tuning. However, given the enormous size of WSIs, it is impractical to perform prompt tuning on all patches. To address this, MSCPT leverages the zero-shot capability of VLMs to preselect high-value patches for subsequent prompt tuning. As shown in Fig. 4a, we conducted an ablation study on the number of patches used for visual prompt tuning on the TCGA-BRCA dataset. The results show that increasing the number of patches for visual prompt tuning improves AUC but significantly increases inference time and GPU memory usage. Therefore, we adopt a trade-off by selecting 30 patches per class, which achieves strong performance while maintaining relatively fast inference time.
Similarly, an ablation study was conducted on the number of prompt sets used for patch selection on the UBC-OCEAN dataset (Fig. 4b). When patches were randomly selected, the model’s performance showed a significant drop, accompanied by an increase in standard deviation. In contrast, increasing the number of prompt sets for patch selection improved the model’s robustness. Moreover, the number of prompt sets had a minimal impact on the results. Considering these factors, we selected 50 prompt sets for patch selection.
Figure 5: Visualization of the original WSI, similarity score map for patch selection, heatmaps generated using MSCPT, attention heatmaps of Metaprompt and the top traditional method, selected high similarity patches and patches with highest similarity scores. The area surrounded by the red line in the WSI indicates the tumor area. The similarity score is derived from the image-text similarity of the NPCGP module.
4.4.3 Effects of Different Prompts in MHPT
We conduct an ablation study on base-to-new generalization using various prompt combinations in Hierarchical Prompt Tuning (MHPT), with the results presented in Table 4. The baseline method utilizes only global prompts. To eliminate the influence of high-level text prompts, we replace the high-level visual descriptions with meaningless placeholders. Similarly, the impact of low-level text prompts is analyzed by removing the prompt generator and excluding low-level text prompts from the MHPT framework. The experimental results demonstrate that both high-level and low-level text prompts positively contribute to the model’s performance. Adding high-level text prompts results in the most significant performance improvement compared to the baseline. Unlike providing pathological prior knowledge to CoOp and Metaprompt, our subsequent ISGPT module fully leverages detailed, fragment-based representations of prior knowledge to enhance performance. When the prompt generator integrates low-level visual descriptions into the prompted text encoder, the model performance further improves. Notably, even with a frozen text encoder, the presence of the prompt generator enables MSCPT to better adapt to low-level visual descriptions.
4.4.4 Effects of Image-text Similarity-based Graph Prompt Tuning
MSCPT utilizes image-text similarity to build a graph neural network that effectively captures the global information in WSIs. To assess the efficacy of the Image-text Similarity-based Graph Prompt Tuning (ISGPT) module, we conducted a comprehensive ablation study. Experimental results on TCGA-RCC are summarized in Table 5.
First, we replaced the ISGPT module with a Transformer, which, in theory, can dynamically learn the contextual relationships between image embeddings and text embeddings. Specifically, we concatenated the dual-scale image and text embeddings obtained after four encoding layers and fed them into a Self-Attention module to dynamically capture the global information of WSIs. However, due to the enormous size of pathology images, directly applying the native Self-Attention mechanism is impractical. To address this, we utilized two alternative approaches: a modified version of TransMIL [1] (excluding the prediction head) and FlashAttention (FA) [44] to perform the experiments. However, experiments show that replacing the ISGPT module with a Transformer resulted in a significant performance drop. This is due to the reparameterization nature of the Transformer, which prevents it from being sufficiently trained in a few-shot setting.
To validate the effectiveness of constructing adjacency matrices via image-text similarity (Sim.), we compared it with k-nearest neighbors (KNN) using patch coordinates (KNN(Coord.)) or visual features (KNN(Feat.)), as in prior studies [26, 27]. Additionally, we tested the effectiveness of GCN by comparing them with GAT [45] and GraphSAGE [46]. Switching to KNN to construct the adjacency matrix led to decreased performance across all metrics, whether based on coordinates or patch visual features. This decline may be attributed to the limited scope of connectivity in the adjacent matrix construction methods. Connecting only nearby patches based on coordinates restricts the GNN to the local context, while connecting visually similar patches based on their features may lack global information about interactions between different types of tissue organization. In contrast, our ISGPT module connects patches related to specific cancer types, overcoming local or visual similarity connection limitations and enabling a more comprehensive contextual understanding. When we replaced GCN with GAT or GraphSAGE, the model’s performance also experienced varying degrees of decline. We demonstrate that complex graph neural networks are unsuitable for few-shot scenarios.
Table 6: Ablation experiment of different instance aggregation methods.
Methods TCGA-NSCLC (PLIP-based, 16-shot)
@@ 4.4.4 Effects of Image-text Similarity-based Graph Prompt Tuning
4.4.5 Effects of Instance Aggregation
To validate the effectiveness of our instance aggregation method, we compared Non-Parametric Cross-Guided Pooling (NPCGP) with other aggregation methods (i.e., Mean, Max, and Attention-based Pooling). The results based on PLIP on the TCGA-NSCLC are presented in Table 6. When we replaced NPCGP with other methods, the performance of the models decreased to varying degrees. This implies that our NPCGP can discern more impactful patches and aggregate them into bag features. The visualization results in Section 4.5 support this point. We also conducted an ablation study on cross-guidance. Instead of computing cross-scale cosine similarity during feature aggregation, we only calculated cosine similarity between image and text embeddings at the same scale. Removing cross-guidance led to a drop in performance across all metrics, likely because LLMs may produce descriptions with incorrect scales. In addition, NPCGP utilizes image-text similarity to classify each WSI, which enhances the interpretability of MSCPT (see Fig. 6 and Section 4.6).
Table 7: Ablation experiment of different large language models.
Methods TCGA-RCC (CLIP-based, 16-shot)
@@ 4.4.5 Effects of Instance Aggregation
Figure 6: Interpretability analysis on the UBC-OCEAN dataset. Including: (a) original WSI, (b) heatmap generated by MSCPT, (c) selected patches with the highest similarity score, (d) description contributions for the selected patch, and (e) description contributions for WSI.
4.4.6 Effects of Large Language Models
@@
To verify the impact of different LLMs on model performance, we compared the performance of MSCPT when using descriptions generated by different LLMs (i.e., Gemini-1.5-pro [47], Claude-3 [48], Llama-3 [49], GPT-3.5 [50], and GPT-4 [43]). The results obtained using CLIP on TCGA-RCC are presented in Table 7. When generating descriptions using Claude-3, MSCPT performs comparably to the baseline. However, MSCPT outperforms the baseline when using other LLMs. This demonstrates MSCPT’s robustness across different LLMs and underscores the importance of accurate visual descriptions in enhancing model performance.
4.5 Visualization Analysis
@@
We performed visualizations of MSCPT, Metaprompt, and the best-performing traditional MIL-based method on one case of PLIP-based TCGA-RCC and one case of CLIP-based TCGA-BRCA. As depicted in Fig. 5a, during patch selection, CLIP assigned high similarity scores to tumor regions as expected, but unexpectedly also to non-tumor areas. This outcome arose because CLIP was not specifically designed for pathological images, resulting in a less-than-optimal zero-shot capability for this type of imagery. However, after prompt tuning using MSCPT, the model correctly assigned high scores to the actual tumor regions, while the regions that originally received high scores dropped to lower score ranges (red arrows). Meanwhile, CLAM-SB struggled to differentiate tumors and non-tumors. Similarly, Metaprompt assigned high attention weights to certain non-tumor tissues (red arrows).
@@
When selecting patches using PLIP, the model could roughly identify tumor regions but also assigned high confidence scores to some non-tumor areas (Fig. 5b). This limitation was effectively addressed by both MSCPT and Metaprompt, as indicated by the red arrows. While ABMIL could also determine instance importance, it tended to assign higher scores to certain non-tumor regions compared to MSCPT and Metaprompt (yellow arrows). Due to PLIP’s improved ability to represent pathological images, Metaprompt produced visualization results comparable to MSCPT. The patches with high similarity scores exhibit features such as enlarged nuclei, increased nuclear density, and irregularity, which are key characteristics of cancer.
4.6 Interpretability Analysis
The probability distribution output by MSCPT is derived from the similarity between the pathological image and the pathological visual description. Therefore, this similarity can be regarded as an indicator of the contribution to the final classification decision. The more an image resembles a specific pathological visual description for a given category, the more likely MSCPT is to classify the WSI as that category. We visualize the similarity between a given pathological image and all pathological visual descriptions within that category, which allows us to assess the contribution of each description to the image’s classification. Similarly, by averaging the similarities selected using Top-K, we can assess each description’s contribution to the overall WSI classification.
As shown in Fig. 6, we performed an interpretability analysis on one case each from the Clear Cell Carcinoma and Endometrioid Carcinoma of the UBC-OCEAN dataset. Panels (d) and (e) show the contribution of pathological visual descriptions to the selected patch and WSI at high resolution, highlighting only the top five and bottom five descriptions. For the selected patch in Clear Cell Carcinoma, the two most contributing pathological visual descriptions are: “Cytoplasmic vacuoles create a honeycomblike pattern in areas with highvacuolization” and “Intracytoplasmic vacuoles push nuclei to the periphery, forming signet-ring-like appearances”. The descriptions corresponding to the selected image from Endometrioid Carcinoma are: “Reactive fibroblasts with elongated nuclei surround tumor nests, contributing to stromal desmoplasia” and “Cytoplasmic eosinophilia creates a pink, granular appearance, contrasting with surrounding clear or vacuolated cells”. These pathological visual descriptions closely match the features and patterns observed in the image, demonstrating a high degree of consistency with its content.
5 Conclusion and Discussion
Conclusion.   In this paper, we propose Multi-Scale and Context-focused Prompt Tuning (MSCPT) to address the Few-shot Weakly-supervised WSI Classification (FSWC) task. MSCPT leverages GPT-4 [43] to generate multi-scale pathological visual descriptions, which guide hierarchical prompt tuning and instance aggregation. Experimental results on five WSI datasets, three downstream tasks, and three Vision-Language Models (VLMs) demonstrate that MSCPT achieves state-of-the-art performance in FSWC tasks. Additionally, MSCPT enables effective fine-tuning of diverse VLMs for WSI-level tasks while offering significant advantages in interpretability compared to traditional MIL-based methods.
Limitations.   However, we observe substantial variations in performance across different datasets and VLMs, primarily due to the dependence of fine-tuning effectiveness on the quality and suitability of pre-trained VLMs. With the rapid advancements in computational pathology, increasingly comprehensive and powerful pre-trained pathological VLMs are emerging, greatly facilitating the resolution of FSWC tasks. We believe that MSCPT, when paired with these advanced VLMs, will serve as a robust and versatile tool in FSWC task and computational pathology.
Future Work.   We outline three key strategic directions for future research: 1) Multimodal Large Language Models: More powerful multimodal large language models (MLLMs) specifically designed for pathology are rapidly advancing [36, 35]. We plan to extend the idea of MSCPT to MLLMs to achieve better performance; 2) Multi-Modal and Multi-Omics Integration: Integrating WSI-derived visual patterns with genomic data is a promising direction. By creating a unified framework for combined histomorphological-molecular classification, we strive to unlock deeper insights into disease mechanisms and improve diagnostic precision; 3) Ultra-Data-Efficient Learning: Extending few-shot capabilities to one/zero-shot scenarios via iterative prompt refinement and self-supervised pretext tasks. Advancing developments in these areas will help bridge the gap between AI and clinical practice, ultimately fostering more interpretable and better performing diagnostic outcomes.